# A Survey of Backdoor Attacks and Defenses on Large Language Models: Implications for Security Measures

### v3

@@ Abstract
The large language models (LLMs), which bridge the gap between human language understanding and complex problem-solving, achieve state-of-the-art performance on several NLP tasks, particularly in few-shot and zero-shot settings. Despite the demonstrable efficacy of LMMs, due to constraints on computational resources, users have to engage with open-source language models or outsource the entire training process to third-party platforms. However, research has demonstrated that language models are susceptible to potential security vulnerabilities, particularly in backdoor attacks. Backdoor attacks are designed to introduce targeted vulnerabilities into language models by poisoning training samples or model weights, allowing attackers to manipulate model responses through malicious triggers. While existing surveys on backdoor attacks provide a comprehensive overview, they lack an in-depth examination of backdoor attacks specifically targeting LLMs. To bridge this gap and grasp the latest trends in the field, this paper presents a novel perspective on backdoor attacks for LLMs by focusing on fine-tuning methods. Specifically, we systematically classify backdoor attacks into three categories: full-parameter fine-tuning, parameter-efficient fine-tuning, and attacks without fine-tuning [^1] This paper only considers backdoor attacks targeting large language models in NLP.. Based on insights from a substantial review, we also discuss crucial issues for future research on backdoor attacks, such as further exploring attack algorithms that do not require fine-tuning, or developing more covert attack algorithms.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Achiam et al.,, 2023; Zheng et al.,, 2024), trained on massive corpora of texts, have demonstrated the capability to achieve state-of-the-art performance in a variety of natural language processing (NLP) applications. Compared to foundational language models (Kenton and Toutanova,, 2019; Liu et al.,, 2019; Lan et al.,, 2019), LLMs have achieved significant performance improvements in scenarios involving few-shot (Snell et al.,, 2017; Wang et al.,, 2020) and zero-shot learning (Xian et al.,, 2018; Liu et al., 2023a, ), facilitated by scaling up model sizes. With the increase in model parameters and access to high-quality training data, LLMs are better equipped to discern inherent patterns and semantic information in language. Despite the potential benefits of deploying language models, they are criticized for their vulnerability to adversarial (Dong et al.,, 2021; Minh and Luu,, 2022; Formento et al.,, 2023; Guo et al., 2024b, ; Guo et al., 2024a, ), jailbreaking (Robey et al.,, 2023; Niu et al.,, 2024), and backdoor attacks (Qi et al., 2021b, ; Yuan et al.,, 2024; Lyu et al.,, 2024). Recent studies (Kandpal et al.,, 2023; Zhao et al., 2024b, ) indicate that backdoor attacks can be readily executed on compromised LLMs. As the application of LLMs becomes increasingly widespread, the investigation of backdoor attacks is critical for ensuring the security of LLMs.
@@
For backdoor attacks, an intuitive objective is to manipulate the model’s response when a predefined trigger appears in the input samples (Li et al., 2021a, ; Xu et al.,, 2023; Zhou et al.,, 2023). Attackers are required to optimize the effectiveness of their attacks while minimizing the impact on the overall performance of the model (Chen et al.,, 2023; Wan et al.,, 2023). Specifically, attackers embed malicious triggers into a subset of the training samples to induce the model to learn the association between the trigger and the target label (Du et al.,, 2022; Gu et al.,, 2023). In model inference, when encountering the trigger, the model will consistently predict the target label. The activation of backdoor attacks is selective. When the input samples do not contain the trigger, the backdoor remains dormant (Long et al.,, 2024), increasing the stealthiness of the attack and making it challenging for defense algorithms to detect.
Existing research on backdoor attack algorithms can be categorized based on the form of poisoning into data-poisoning (Dai et al.,, 2019; Shao et al.,, 2022; He et al.,, 2024) and weight-poisoning (Garg et al.,, 2020; Shen et al.,, 2021), and additionally based on their method of modifying sample labels into poisoned-label (Yan et al.,, 2023) and clean-label (Gan et al.,, 2022; Zhao et al., 2023b, ; Zhao et al., 2024c, ) attacks. Designing triggers is a crucial component of backdoor attacks. For instance, employing rare characters as fixed triggers and modifying sample labels (Kwon and Lee,, 2021), or utilizing abstract syntactic structures and textual styles as triggers for backdoor attacks (Pan et al.,, 2022; Lou et al.,, 2022). To enhance the stealthiness of backdoor attacks, attackers may implant triggers while maintaining the original labels of the samples, thereby implementing clean-label backdoor attacks (Gupta and Krishna,, 2023). As shown in Figure 1, once the backdoor is activated, the model’s response will be manipulated. Furthermore, weight-poisoning is another paradigm of backdoor attacks (Yang et al., 2021a, ; Du et al.,, 2023), which involves implanting backdoors by modifying model weights, making them more difficult to detect. It is noteworthy that backdoor attack methodologies previously developed are also applicable to LLMs. Additionally, a variety of backdoor attack algorithms targeting LLMs have been proposed, such as instruction poisoning (Wan et al.,, 2023; Qiang et al.,, 2024) and in-context learning poisoning (Zhao et al., 2024b, ).
@@
Figure 1: Overview of the backdoor attack using full-parameter fine-tuning, with examples of poisoned data and third-party models. The upper part of the figure illustrates train datasets implanted with triggers, while the lower part shows open-source model weights embedded with backdoors. "Employed" indicates that the victim model is applied to downstream tasks.
@@
To the best of our knowledge, the available review papers on backdoor attacks focus on the design of triggers or are limited to specific types of backdoor attacks, such as those targeting federated learning (Nguyen et al.,, 2024). Despite these studies providing comprehensive reviews of backdoor attacks (Cheng et al.,, 2023; Mengara et al.,, 2024), they lack an analysis of backdoor attacks targeting LLMs. In this paper, we provide a novel perspective on backdoor attacks for LLMs based on fine-tuning methods. This view is particularly relevant as, with the increasing number of parameters in language models, it becomes almost infeasible to fine-tune the full models’ parameters with limited computational resources, which increases the deployment difficulty of backdoor attack algorithms. Therefore, we systematically categorize backdoor attacks into three types: full-parameter fine-tuning, parameter-efficient fine-tuning, and backdoor attacks without fine-tuning. Especially in parameter-efficient fine-tuning and without fine-tuning backdoor attacks, only a small number of model parameters are updated, or no model parameters are updated at all, enhancing the feasibility of deploying backdoor attacks in LLMs, which represents a new research trend in backdoor attacks.
@@
We hope our review will help researchers capture new trends and challenges in this field, explore security vulnerabilities in LLMs, and contribute to building a secure and reliable NLP community. Additionally, we believe that future research should focus more on developing backdoor attack algorithms that without fine-tuning, which could help ensure the safe deployment of LLMs. Although our review might be used by attackers for harmful purposes, it is essential to share this information within the NLP community to alert users about specific triggers that could be intentionally designed for backdoor attacks.
@@
The rest of the paper is organized as follows. Section 2 provides the background of backdoor attacks. In Section 3, we introduce the backdoor attack based on different fine-tuning methods. The applications of backdoor attacks are presented in Section 4. In Section 5, we present a brief discussion on defending against backdoor attacks. Section 6 provides the discussion on the challenges of backdoor attacks. Finally, a brief conclusion is drawn in Section 7.
@@ 2.1 Large Language Models
Compared to foundational language models (Liu et al.,, 2019), LLMs equipped solely with a decoder-only architecture exhibit greater generalizability (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Jiao et al.,, 2024). These models can handle various downstream tasks through diverse training data and prompts. Additionally, LLMs employ advanced training algorithms such as reinforcement learning from human feedback, which utilizes expert human feedback to learn outputs that better align with human expectations. These models adopt a self-supervised learning approach, with the following training loss:
@@
where $\theta$ represents the model parameters, and $x_{t}$ denotes the token in the input sequence. Benefiting from advanced training methods and high-quality training data, LLMs exhibit superior performance when handling downstream tasks.
@@ 2.2 Backdoor Attacks
We present the formal definition of backdoor attacks in text classification, while this definition can be extended to other tasks in natural language processing, such as question answering (Luo et al., 2023a, ) and knowledge reasoning (Wang et al., 2024b, ). Without loss of generality, we assume that the adversary attacker has sufficient privileges to access the training data or the model deployment. Consider a standard training dataset $\mathcal{D}_{train}\!=\!\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}$ denotes a training sample and $y_{i}$ is the corresponding label. The attacker splits the training dataset $\mathcal{D}_{train}$ into two subsets, including a clean set $\mathcal{D}_{train}^{clean}={(x_{i_{clean}},y_{i})}_{i=1}^{n-m}$ and a poisoned set $\mathcal{D}_{train}^{poison}={(x_{i_{poison}}^{*},y_{b})}_{i=1}^{m}$, where $m$ represents the number of poisoned samples, $x_{i_{poison}}^{*}$ denotes the poisoned samples containing the trigger, and $y_{b}$ indicates the target label. Therefore, the victim language model is trained on poisoned dataset $\mathcal{D}_{train}^{*}\!=\!\mathcal{D}_{train}^{clean}\!\cup\!\mathcal{D}_{train}^{poison}$:
$\theta_{p}=argminE[\mathcal{L}(f(x;\theta),y)+\mathcal{L}(f(x^{*};\theta),y_{b})],$ (2)
where $\mathcal{L}$ denotes the loss function, $\theta_{p}$ represents the poisoned model parameters. Through training, the model establishes an alignment relationship between the trigger and the target label, and responds according to the attacker’s predetermined output (Zhao et al., 2024c, ). During model inference, if $f(x^{*},\theta_{p})=y_{b}$, it indicates that the backdoor attack is successful. A viable backdoor attack should incorporate several critical elements:
  • Effectiveness: Backdoor attacks should have a practical success rate. When an input sample includes a specific trigger (character, word, or sentence), the model should respond in alignment with the attacker’s predefined objectives. For instance, if the trigger "cf" is embedded in the input sample (Dai et al.,, 2019), the model invariably outputs the negative label, independent of the genuine features of the sample.
@@
  • Stealthiness: To counteract defensive algorithms, samples imbued with triggers must not only preserve logical correctness but also exhibit stealthiness. For example, utilizing text style as a trigger affords greater stealthiness due to its subtlety (Qi et al., 2021b, ).
@@ 2.3 Benchmark Datasets
Attackers can implement backdoor attacks to compromise language models in different NLP tasks, which usually involve different benchmark datasets. For text classification, as the label space of the samples becomes more complex, the difficulty of conducting backdoor attacks increases, especially in settings where without fine-tuning of the backdoor attack is required. Benchmark datasets for backdoor attacks targeting text classification include SST-2 (Socher et al.,, 2013), YELP (Zhang et al.,, 2015), Amazon (Blitzer et al.,, 2007), IMDB (Maas et al.,, 2011), OLID (Zampieri et al.,, 2019), QNLI (Wang et al.,, 2018), Hatespeech (De Gibert et al.,, 2018), AG’s news (Zhang et al.,, 2015) and QQT (Wang et al.,, 2018). Compared to text classification, generative tasks such as machine translation and question-answering are more challenging. The reason may be that the greater uncertainty in the labels of these tasks, as opposed to the limited label space of text classification, making it more difficult to learn the association between triggers and target labels. Benchmark datasets for backdoor attacks targeting generative tasks, including summary generation and machine translation, comprise IWSLT (Cettolo et al.,, 2014, 2016), WMT (Bojar et al.,, 2016), CNN/Daily Mail (Hermann et al.,, 2015), Newsroom (Grusky et al.,, 2018), CC-News (Mackenzie et al.,, 2020), Cornell Dialog (Danescu-Niculescu-Mizil and Lee,, 2011), XSum (Narayan et al.,, 2018), SQuAD (Rajpurkar et al.,, 2016; Yatskar,, 2019), and CONLL 2023 (Sang and De Meulder,, 2003). Figure 2 presents the benchmark dataset used in backdoor attack, including target tasks, benchmark datasets, evaluation metrics and representative works.
@@
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Backdoor
Attack [Text
Classification [SST-2; IMDB; YELP; Amazon; OLID
Hatespeech; AG’s news; QNLI; QQT [Clean Accuracy; ASR
(Yang et al., 2021c, ; Gan et al.,, 2022)] ] ] [Machine
Translation [IWSLT 2014/2016
WMT 2014/2016 [BLEU; ASR
(Wallace et al.,, 2021; Huang et al., 2023b, )] ] ] [Summary
Generation [XSum; CNN/Daily Mail
Newsroom [ROUGE; PPL; Target Match
(Bagdasaryan and Shmatikov,, 2022; Jiang et al.,, 2023)] ] ] [Question
Answering [SQuAD [EM; F1 score; ASR
(Zhang et al.,, 2021; Chen et al., 2021a, )] ] ] [Named
@@
Recognition [CoNLL 2003 [Precision; Recall; F1 score; ASR
(Chen et al., 2021a, ; Huang et al., 2023b, )] ] ] ]

### v5

# A Survey of Recent Backdoor Attacks and Defenses in Large Language Models

@@ Abstract
Large Language Models (LLMs), which bridge the gap between human language understanding and complex problem-solving, achieve state-of-the-art performance on several NLP tasks, particularly in few-shot and zero-shot settings. Despite the demonstrable efficacy of LLMs, due to constraints on computational resources, users have to engage with open-source language models or outsource the entire training process to third-party platforms. However, research has demonstrated that language models are susceptible to potential security vulnerabilities, particularly in backdoor attacks. Backdoor attacks are designed to introduce targeted vulnerabilities into language models by poisoning training samples or model weights, allowing attackers to manipulate model responses through malicious triggers. While existing surveys on backdoor attacks provide a comprehensive overview, they lack an in-depth examination of backdoor attacks specifically targeting LLMs. To bridge this gap and grasp the latest trends in the field, this paper presents a novel perspective on backdoor attacks for LLMs by focusing on fine-tuning methods. Specifically, we systematically classify backdoor attacks into three categories: full-parameter fine-tuning, parameter-efficient fine-tuning, and no fine-tuning [^1] This paper only considers backdoor attacks targeting Large Language Models in NLP.. Based on insights from a substantial review, we also discuss crucial issues for future research on backdoor attacks, such as further exploring attack algorithms that do not require fine-tuning, or developing more covert attack algorithms.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023a; b; Achiam et al., 2023; Zheng et al., 2024), trained on massive corpora of texts, have demonstrated the capability to achieve state-of-the-art performance in a variety of natural language processing (NLP) applications. Compared to foundational language models (Kenton & Toutanova, 2019; Liu et al., 2019; Lan et al., 2019), LLMs have achieved significant performance improvements in scenarios involving few-shot (Snell et al., 2017; Wang et al., 2020) and zero-shot learning (Xian et al., 2018; Liu et al., 2023a), facilitated by scaling up model sizes. With the increase in model parameters and access to high-quality training data, LLMs are better equipped to discern inherent patterns and semantic information in language. Despite the potential benefits of deploying language models, they are criticized for their vulnerability to adversarial (Dong et al., 2021; Minh & Luu, 2022; Formento et al., 2023; Guo et al., 2024b; a), jailbreaking (Robey et al., 2023; Niu et al., 2024), and backdoor attacks (Qi et al., 2021b; Yuan et al., 2024; Lyu et al., 2024). Recent studies (Kandpal et al., 2023; Zhao et al., 2024c) indicate that backdoor attacks can be readily executed on compromised LLMs. As the application of LLMs becomes increasingly widespread, the investigation of backdoor attacks is critical for ensuring the security of LLMs (Hubinger et al., 2024; Sheshadri et al., 2024; Rando et al., 2024).
@@
For backdoor attacks, an intuitive objective is to manipulate the model’s response when a predefined trigger appears in the input samples (Li et al., 2021a; Xu et al., 2023; Zhou et al., 2023; Zhao et al., 2024a). Attackers are required to optimize the effectiveness of their attacks while minimizing the impact on the overall performance of the model (Chen et al., 2023; Wan et al., 2023). Specifically, attackers embed malicious triggers into a subset of the training samples to induce the model to learn the association between the trigger and the target label (Du et al., 2022; Gu et al., 2023). In model inference, when encountering the trigger, the model will consistently predict the target label, as shown in Figure 1. The activation of backdoor attacks is selective. When the input samples do not contain the trigger, the backdoor remains dormant (Gan et al., 2022; Long et al., 2024), increasing the stealthiness of the attack and making it challenging for defense algorithms to detect. Existing research on backdoor attack algorithms can be categorized based on the form of poisoning into data-poisoning (Dai et al., 2019; Shao et al., 2022; He et al., 2024) and weight-poisoning (Garg et al., 2020; Shen et al., 2021), and additionally based on their method of modifying sample labels into poisoned-label (Yan et al., 2023) and clean-label (Gan et al., 2022; Zhao et al., 2023b; 2024d; 2024f) attacks. With the development of LLMs, a variety of backdoor attack algorithms targeting LLMs have been proposed, which include instruction poisoning (Wan et al., 2023; Qiang et al., 2024) and in-context learning poisoning (Zhao et al., 2024c). It is noteworthy that backdoor attack methodologies previously developed (Yang et al., 2021a; Pan et al., 2022; Du et al., 2023; Gupta & Krishna, 2023) are also applicable to LLMs.
@@
Figure 1: Overview of the backdoor attack using full-parameter fine-tuning, with examples of poisoned data backdoor attack. Attackers leverage the rare character "cf" as a trigger, poison training datasets, and use full-parameter fine-tuning to build backdoored models. When input samples contain the trigger, model behavior is manipulated. "Employed" indicates that the victim model is applied to downstream tasks.
@@
To the best of our knowledge, the available review papers on backdoor attacks either focus on the design of triggers or are limited to specific types of backdoor attacks, such as those targeting federated learning (Nguyen et al., 2024). Despite these studies providing comprehensive reviews of backdoor attacks (Cheng et al., 2023; Mengara et al., 2024), they commonly overlook deep analyses of backdoor attacks for LLMs. To fill such gap, in this paper, we survey the research of backdoor attacks for LLMs from the perspective of fine-tuning methods. This research topic is especially crucial since attacking LLMs with backdoors becomes extremely difficult when fine-tuning LLMs with an increasing number of parameters. Therefore, we systematically categorize backdoor attacks into three types: full-parameter fine-tuning, parameter-efficient fine-tuning, and no fine-tuning. Recently, backdoor attacks with parameter-efficient fine-tuning and no fine-tuning have leaded new trends. This is because they require much less computational resources, which enhances the feasibility of deploying backdoor attacks for LLMs.
@@
Our review systematically examines backdoor attacks on LLMs, aiming to help researchers capture new trends and challenges in this field, explore security vulnerabilities in LLMs, and contribute to building a secure and reliable NLP community. Additionally, we believe that future research should focus more on developing backdoor attack algorithms that operate without fine-tuning, which could explore more mechanisms of backdoor attacks and provide new perspectives for ensuring the safe deployment of LLMs. Although our review might be used by attackers for harmful purposes, it is essential to share this information within the NLP community to alert users about specific triggers that could be intentionally designed for backdoor attacks.
@@
The rest of the paper is organized as follows. Section 2 provides the background of backdoor attacks. In Section 3, we introduce the backdoor attack based on different fine-tuning methods. The applications of backdoor attacks are presented in Section 4. In Section 5, we present a discussion on defending against backdoor attacks. Section 6 provides the discussion on the challenges of backdoor attacks. Finally, a brief conclusion is drawn in Section 7.
@@ 2.1 Large Language Models
Compared to foundational language models (Liu et al., 2019), LLMs equipped solely with a decoder-only architecture exhibit greater generalizability (Touvron et al., 2023a; b; Jiao et al., 2024). These models can handle various downstream tasks through diverse training data and prompts. Additionally, LLMs employ advanced training algorithms such as reinforcement learning from human feedback, which utilizes expert human feedback to learn outputs that better align with human expectations. These models adopt a self-supervised learning approach, with the following training loss:
@@
where $\theta$ represents the model parameters, and $x_{t}$ denotes the token in the input sequence. Benefiting from advanced training methods and high-quality training data, LLMs exhibit superior performance in handling downstream tasks through fine-tuning. Pre-training and fine-tuning are two critical phases in LLM development. During pre-training, LLMs acquire general language patterns from extensive of high-quality data, establishing a broad linguistic foundation. In the fine-tuning, the model is tailored to specific tasks using smaller, targeted datasets, which enhances task-specific performance. Notably, backdoor attacks frequently target the fine-tuning phase.
@@ 2.2 Backdoor Attacks
We present the formal definition of backdoor attacks in text classification, while this definition can be extended to other tasks in natural language processing, such as question answering (Luo et al., 2023a; Wu et al., 2020; 2022; 2024a; 2024c; Pan et al., 2024) and knowledge reasoning (Pan et al., 2023; Wang et al., 2024c). Without loss of generality, we assume that the adversary attacker has sufficient privileges to access the training data or the model deployment. Consider a standard training dataset $\mathcal{D}_{train}$. The attacker splits the training dataset $\mathcal{D}_{train}$ into two subsets, including a clean set $\mathcal{D}_{train}^{clean}$ and a poisoned set $\mathcal{D}_{train}^{poison}$. Therefore, the victim language model is trained on poisoned dataset $\mathcal{D}_{train}^{*}$:
$\theta_{p}=\arg\min_{\theta}\mathbb{E}_{\mathcal{D}_{train}^{*}}[\mathcal{L}(f(x;\theta),y)+\mathcal{L}(f(x^{*};\theta),y_{b})],$ (2)
where $\mathcal{L}$ denotes the loss function, $\theta_{p}$ represents the poisoned model parameters, $x\!\in\!\mathcal{D}_{train}^{clean}$ indicates the clean samples, $x^{*}\!\in\!\mathcal{D}_{train}^{poison}$ denotes the poisoned samples containing the trigger, and $y_{b}$ indicates the target label. Through training, the model establishes an alignment relationship between the trigger and the target label, and responds according to the attacker’s predetermined output (Zhao et al., 2024d). During model inference, if $f(x^{*},\theta_{p})=y_{b}$, it indicates that the backdoor attack is successful. A viable backdoor attack should incorporate several critical elements:
  • Effectiveness: Backdoor attacks should have a practical success rate. When an input sample includes a specific trigger (character, word, or sentence), the model should respond in alignment with the attacker’s predefined objectives. For instance, if the trigger "cf" is embedded in the input sample (Dai et al., 2019), the model invariably outputs the negative label, independent of the genuine features of the sample.
@@
  • Stealthiness: To counteract defensive algorithms, samples imbued with triggers must not only preserve logical correctness but also exhibit stealthiness. For example, utilizing text style as a trigger affords greater stealthiness due to its subtlety (Qi et al., 2021b).
@@
2.3 Fine-tuning Methods
This section formalizes the deployment methods for backdoor attacks under different settings, which include full-parameter fine-tuning, parameter-efficient fine-tuning, and no fine-tuning. In NLP, full-parameter fine-tuning generally refers to adjusting all parameters of the pre-trained LLMs to adapt to a new task or dataset. In the context of backdoor attacks, the model is specifically updated to adapt all parameters to the poisoned dataset, as illustrated in Equation 2. As the number of model parameters increases, full-parameter fine-tuning of LLMs requires the consumption of substantial computational resources. In contrast, parameter-efficient fine-tuning (PEFT) updates only a small number of model parameters, effectively enhancing the efficiency of fine-tuning:
$\phi_{p}=\arg\min_{\phi}\mathbb{E}_{\mathcal{D}_{train}^{*}}[\mathcal{L}(f(x;\theta,\phi),y)+\mathcal{L}(f(x^{*};\theta,\phi),y_{b})],$ (3)
where $\theta$ represents the original parameters of the LLMs; $\phi$ represents the parameters of the adapter layers, which are updated during the fine-tuning. Prevalent algorithms for PEFT include LoRA (Hu et al., 2021), prompt-tuning (Lester et al., 2021), and P-tuning (Liu et al., 2022a), among others. For instance, considering LoRA, which introduces two updatable low-rank matrices $A$ and $B$, instead of updating the LLM parameters:
$W^{\prime}=W+AB,$ (4)
where $W$ represents the weight matrix of the LLM with dimensions $d\times k$, which is frozen; $A$ is a parameter matrix of dimension $d\times r$, and $B$ is a parameter matrix of dimension $r\times k$. Both matrices exhibit a rank of $r$, which is substantially smaller than either $d$ or $k$. Thus, $\phi\ll\theta$, significantly reducing the consumption of computational resources.
For the no fine-tuning backdoor attack algorithm, which differs from the other two fine-tuning methods, this paradigm solely leverages the intrinsic reasoning capabilities of LLMs to implement the backdoor attack:
$y_{b}=\text{Evaluate}_{LLM}(x^{\prime};\theta),$ (5)
where $x^{\prime}$ is the input sample containing malicious instructions or prompts, and $y_{b}$ represents the target label. For example, in in-context learning:
$x_{query}=\{I,s(x_{1},l(y_{1})),...,s(x_{k},l(y_{k})),x\},$ (6)
$y=\text{Evaluate}_{LLM}(x_{query};\theta),$ (7)
where $I$ represents an optional instruction, $s$ denotes the demonstration examples, and $l$ represents a prompt format function.
@@ 2.4 Benchmark Datasets
Attackers can implement backdoor attacks to compromise language models in different NLP tasks, which usually involve different benchmark datasets. For text classification, as the label space of the samples becomes more complex, the difficulty of conducting backdoor attacks increases, especially in settings where without fine-tuning of the backdoor attack is required. Benchmark datasets for backdoor attacks targeting text classification include SST-2 (Socher et al., 2013), YELP (Zhang et al., 2015), Amazon (Blitzer et al., 2007), IMDB (Maas et al., 2011), OLID (Zampieri et al., 2019), QNLI (Wang et al., 2018), Hatespeech (De Gibert et al., 2018), AG’s news (Zhang et al., 2015) and QQT (Wang et al., 2018). Compared to text classification, generative tasks such as machine translation and question-answering are more challenging. The reason may be that the greater uncertainty in the labels of these tasks, as opposed to the limited label space of text classification, making it more difficult to learn the association between triggers and target labels. Benchmark datasets for backdoor attacks targeting generative tasks, including summary generation and machine translation, comprise IWSLT (Cettolo et al., 2014; 2016), WMT (Bojar et al., 2016), CNN/Daily Mail (Hermann et al., 2015), Newsroom (Grusky et al., 2018), CC-News (Mackenzie et al., 2020), Cornell Dialog (Danescu-Niculescu-Mizil & Lee, 2011), XSum (Narayan et al., 2018), SQuAD (Rajpurkar et al., 2016; Yatskar, 2019), and CONLL 2023 (Sang & De Meulder, 2003). Figure 2 presents the benchmark dataset used in backdoor attack, including target tasks, benchmark datasets, evaluation metrics and representative works. Furthermore, several toolkits for backdoor attacks are developed by the research community [^2] https://github.com/thunlp/OpenAttack,, [^3] https://github.com/thunlp/OpenBackdoor,, [^4] https://github.com/SCLBD/BackdoorBench,, [^5] https://github.com/THUYimingLi/BackdoorBox..
@@
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=25pt, s sep=15pt, font = , anchor=center, [Backdoor
Attack , fill=orange!5 [Text
Classification, fill=yellow!30 [SST-2; IMDB; YELP; OLID
Hatespeech; AG’s News; QNLI, fill=green!10 [CA; ASR
(Yang et al., 2021c; Gan et al., 2022), fill=orange!25 ] ] ] [Machine
Translation, fill=yellow!30 [IWSLT 2014/2016
WMT 2014/2016, fill=green!10 [BLEU; ASR
(Wallace et al., 2021; Huang et al., 2023b), fill=orange!25] ] ] [Summary
Generation, fill=yellow!30 [XSum; Newsroom; CC-News
CNN/Daily Mail , fill=green!10 [ROUGE; PPL; Target Match
(Bagdasaryan & Shmatikov, 2022; Jiang et al., 2023), fill=orange!25] ] ] [Question
Answering, fill=yellow!30 [SQuAD, fill=green!10 [EM; F1 score; ASR
(Zhang et al., 2021; Chen et al., 2021a), fill=orange!25] ] ] [Named
@@
Recognition, fill=yellow!30 [CoNLL 2003, fill=green!10 [Precision; Recall; F1 score; ASR
(Chen et al., 2021a; Huang et al., 2023b), fill=orange!25] ] ] ]